# Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection

### v1

@@ Appendix B Preliminary Terms for Self-Containment
  • Conformal $p$-values. The notion of conformal $p$-values was originally introduced by Vovk et al. [2005] for constructing prediction intervals. A conformal $p$-value quantifies how well a new observation conforms to a reference set, based on a chosen nonconformity score function. More recently, several works have applied conformal $p$-values to sample selection from a multiple testing perspective [Jin and Candès, 2023a, Bates et al., 2023, Wang et al., 2024]. The conformal $p$-values are defined as
@@
Appendix C Additional Details of Our Algorithms
In this Section, we provide additional details of our algorithms, including the concrete testing levels of proposed algorithms in Appendix C.1, Optimized OCTF by adaptively tuned online learning in Appendix C.2, and the additional explanation about utilizing feedback in SAFFRON-type algorithms is in Appendix C.3.
C.1 Testing levels of the proposed approaches
@@
    Testing levels for $\text{LF}_{\text{dep}}$ and $\text{SF}_{\text{dep}}$:
@@
    $\alpha_{t}^{\text{LF}_{\text{dep}}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-r_{1}}\mathbb{I}\{r_{1}<t\}+\alpha\sum_{j=2}^{\infty}\gamma_{t-r_{j}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j},$
    $\alpha_{t}^{\text{SF}_{\text{dep}}}:=\min\left\{\lambda,(1-\lambda)\left(w_{0}\gamma_{t-C_{0+}}+(\alpha-w_{0})\gamma_{t-r_{1}-C_{1+}}+\alpha(\sum_{j\geq 2}\gamma_{t-r_{j}-C_{j+}})\right)+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j}\right\},$
@@
    Testing levels for LFS and SFS: Recall that the proposed LF update rule is given by:
    $\alpha_{t}^{\text{LF}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j},$ (24)
@@
    ${\alpha}_{t}^{\text{LFS}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tilde{\tau}_{1}}\mathbb{I}\{{\tilde{\tau}}_{1}<t\}+\alpha\sum_{j:\tilde{\tau}_{j}<t,\tilde{\tau}_{j}\neq\tilde{\tau}_{1}}\gamma_{t-\tilde{\tau}_{j}}+\sum_{j:j<t}\gamma_{t-j}{\alpha}_{j}\theta_{j},$ (25)
@@
    where $\tilde{\alpha}_{t}^{\text{SFS}}=s_{0}\gamma_{t-C_{0+}}+((1-\lambda)\alpha-s_{0})\gamma_{t-\tilde{\tau}_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tilde{\tau}_{j}-C_{j+}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j}.$
C.2 Optimized OCTF by adaptively tuned online learning
Our framework also allows online learning for the predictive models and the model selection is specific to tuning the parameters of model with online updating. Suppose we have $K$ candidate hyper-parameters. We will implement $K$ corresponding models in parallel throughout our procedure. At each time $t$, the models will be updated by training on the current data $\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{t-1}\cup\{({\bf X}_{t},\tilde{Y}_{t})\}$. Here as $Y_{t}$ is unobserved currently, we will regard $\tilde{Y}_{t}$ as the null value. To select a best model for our testing procedure, we can use a similar criterion as ${\mathcal{M}}^{\mathrm{EWMA}}_{t}(k,{\mathcal{D}}_{t})$. And the conformal p-value is constructed by the model $\tilde{\mu}_{\hat{k}_{t}}^{(t)}$. The detailed pseudo-codes of the adaptively tuned online learning procedure is shown in Algorithm 4.
Algorithm 4 Optimized OCTF by adaptively tuned online learning
0:  Initial data ${\mathcal{D}}_{1}=\{({\bf X}_{i},Y_{i})\}_{i=-N}^{0}$, target region $\mathcal{A}$, FDR target level $\alpha\in(0,1)$, $K$ candidate hyper-parameters $\{\eta_{k}\}_{k=1}^{K}$, loss function $L(\beta,Y;\eta_{k})$ for the predictive model using hyperparameter $\eta_{k}$, evaluating criterion $\mathcal{M}$, parameter $s_{0}$, parameter sequence $\{\gamma_{t}\}$, stopping time $T$.
1:  for $t=1,\dots,T$ do
2:   Observe test data ${\bf X}_{t}$
3:   Online update one-step-ahead models for $k=1,\cdots,K$
[…]
and the corresponding predictive model is $\tilde{\mu}_{k}^{(t)}$. The score $V(\cdot;\tilde{\mu}_{k}^{(t)})$ is determined by the predictive model $\tilde{\mu}_{k}^{(t)}$.
4:   Decide the predictive model for $t$-th test sample by
$\widehat{k}_{t}=\underset{k\in[K]}{\arg\max}\;{\mathcal{M}}^{\text{EWMA}}(k,{\mathcal{D}}_{t})$
as the score function $V(\cdot;\tilde{\mu}_{k}^{(t)})$ is decided by $\tilde{\mu}_{k}^{(t)}$.
5:   Construct optimized conformal p-value by
$p^{\text{opt}}_{t}=\frac{1+\sum_{i\in{\mathcal{C}}_{0t}}\mathbb{I}\{V({\bf X}_{i};\tilde{\mu}_{\hat{k}_{t}}^{(t)})\leq V({\bf X}_{t};\tilde{\mu}_{\hat{k}_{t}}^{(t)})\}}{1+|{\mathcal{C}}_{0t}|}$
6:   Update $\alpha_{t}={\alpha}_{t}^{\text{LFS}}$ in Equation (25) or $\alpha_{t}={\alpha}_{t}^{\text{SFS}}$ in Equation (26)
7:   Make a decision $\delta_{t}=\mathbb{I}\{p_{t}^{\text{opt}}\leq\alpha_{t}\}$
8:   Obtain the revealed feedback $Y_{t}$ (and thus obtain $\theta_{t}$)
9:   Update the calibration dataset ${\mathcal{C}}_{0t}$.
10:  end for
10:  Rejection set $\mathcal{R}_{\text{opt}}=\{i:\delta_{i}=1,~\delta_{i}\in{\bm{\delta}}^{T}\}$.
C.3 Explanation about utilizing feedback in SAFFRON
@@
Here, we explain why this form is not adopted in the proposed GAIF. In practice, we typically have $D_{j}^{+}<C_{j}^{+}$, since the set filtered by $p$-values is larger than the true alternative set. Consequently, due to the exponential decay design of $\gamma_{t}$, we have
@@
Appendix D Extensions of GAIF based on $e$-values
@@
Similar to the GAIF framework, we propose the estimators $\widehat{\operatorname{FDP}}_{\text{e-LF}}$ and $\widehat{\operatorname{FDP}}_{\text{e-SF}}$ as follows:
$\operatorname{FDP}^{*}(t)=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\leq\widehat{\operatorname{FDP}}_{\text{e-LF}}=\sum_{j\leq t-1}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}+\frac{\alpha_{t}}{1+R({t-1})}.$
$\widehat{\operatorname{FDP}}(t)_{\text{e-SF}}=\sum_{j\leq t-1}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}\frac{\mathbb{I}\{e_{j}\leq 1/\lambda_{j}\}}{1-\lambda_{j}}+\frac{\alpha_{t}}{1+R({t-1})}\frac{\mathbb{I}\{e_{t}\leq 1/\lambda_{t}\}}{1-\lambda_{t}}.$
Then we require $\widehat{\operatorname{FDP}}_{\text{e-LF}}\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\text{e-SF}}\leq\alpha$ when constructing the testing levels. The corresponding testing levels are as follows:
$\alpha_{t}^{\text{e-LF}}=\omega_{t}\left(\alpha-\sum_{j=1}^{t-1}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}\right)(R({t-1})+1),$ (28)
$\alpha_{t}^{\text{e-SF}}=\omega_{t}\left(\alpha(1-\lambda)-\sum_{j=1}^{t-1}\frac{\alpha_{j}(1-\theta_{j})\mathbb{I}\{e_{j}<1/\lambda\}}{1+R({j-1})}\right)(R({t-1})+1),$ (29)
@@
The Generalized Alpha-Investing with Feedback procedure based on $e$-values (e-GAIF) is summarized in Algorithm 5. The $e$-LF and $e$-SF can control online FDR validly if the null $e$-values satisfy conditionally valid.
Algorithm 5 $e$-GAIF ($e$-LF and $e$-SF)
@@
3:   Update $\alpha_{t}=\alpha_{t}^{\text{e-LF}}$ in Eq.(28) (or $\alpha_{t}=\alpha_{t}^{\text{e-SF}}$ in Eq.(29))
@@
Theorem 4 (Online FDR control for e-GAIF).
@@
where $\mathcal{F}_{t-1}=\sigma(\delta_{1},\dots,\delta_{t-1};\theta_{1},\dots,\theta_{t-1})$ is the sigma field generated from past rejections and feedback, then if the parameters $\{\alpha_{t}\}_{t\in\mathbb{N}}$ are selected such that $\widehat{\operatorname{FDP}}_{\operatorname{e-LF}}(t)\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\operatorname{e-SF}}(t)\leq\alpha$ , then we have

### v2

@@ A.1 Preliminary terms for self-containment
  • Conformal $p$-values. The notion of conformal $p$-values was originally introduced by Vovk et al. [2005] for constructing prediction intervals. A conformal $p$-value quantifies how well a new observation conforms to a reference set, based on a chosen nonconformity score function. More recently, several works have applied conformal $p$-values to sample selection from a multiple testing perspective [Jin and Candès, 2025, Bates et al., 2023, Wang et al., 2024]. The conformal $p$-values are defined as
@@
A.2 Testing levels of the proposed approaches
@@
    Testing levels for $\text{GAIF}_{\text{dep}}$ and $\text{Ada-GAIF}_{\text{dep}}$:
@@
    $\alpha_{t}^{\text{GAIF}_{\text{dep}}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-r_{1}}\mathbb{I}\{r_{1}<t\}+\alpha\sum_{j=2}^{\infty}\gamma_{t-r_{j}}+\sum_{j:j\in\mathcal{I}_{t}}\gamma_{j}\alpha_{j}\theta_{j},$
    $\alpha_{t}^{\text{Ada-GAIF}_{\text{dep}}}:=\min\left\{\lambda,(1-\lambda)\left(w_{0}\gamma_{t-C_{0+}}+(\alpha-w_{0})\gamma_{t-r_{1}-C_{1+}}+\alpha(\sum_{j\geq 2}\gamma_{t-r_{j}-C_{j+}})\right)+\sum_{j:j\in\mathcal{I}_{t}}\gamma_{j}\alpha_{j}\theta_{j}\right\},$
@@
    Testing levels for LFS and SFS: Recall that the proposed test levels of GAIF under full and instant feedback, i.e., $\alpha_{t}^{\rm LF}$, are given by:
    $\alpha_{t}^{\text{LF}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}}+\sum_{j:j<t}\gamma_{j}\alpha_{j}\theta_{j},$ (17)
@@
    ${\alpha}_{t}^{\text{LFS}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tilde{\tau}_{1}}\mathbb{I}\{{\tilde{\tau}}_{1}<t\}+\alpha\sum_{j:\tilde{\tau}_{j}<t,\tilde{\tau}_{j}\neq\tilde{\tau}_{1}}\gamma_{t-\tilde{\tau}_{j}}+\sum_{j:j<t}\gamma_{j}{\alpha}_{j}\theta_{j},$ (18)
@@
    where $\tilde{\alpha}_{t}^{\text{SFS}}=s_{0}\gamma_{t-C_{0+}}+((1-\lambda)\alpha-s_{0})\gamma_{t-\tilde{\tau}_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tilde{\tau}_{j}-C_{j+}}+\sum_{j:j<t}\gamma_{j}\alpha_{j}\theta_{j}.$
A.3 Choice of non-conformity score function $V$
Appendix B Additional Discussions on GAIF and Adaptive GAIF
In this section, we provide additional details of our GAIF and Adaptive GAIF algorithms in Sections 2 and 3.
B.1 Explanation about utilizing feedback in SAFFRON
@@
Here, we explain why this form is not adopted in the proposed GAIF. In practice, we typically have $D_{j}^{+}<C_{j}^{+}$, since the set filtered by $p$-values is larger than the true non-null set. Consequently, due to the exponential decay design of $\gamma_{t}$, we have
@@
B.2 Extensions of GAIF based on $e$-values
@@
Similar to the GAIF framework, we propose the estimates $\widehat{\operatorname{FDP}}_{\text{e-GAIF}}$ and $\widehat{\operatorname{FDP}}_{\text{e-Ada-GAIF}}$ as follows:
$\operatorname{FDP}^{*}(t)=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\leq\widehat{\operatorname{FDP}}_{\text{e-GAIF}}=\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}+\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}}{1+R({t-1})}.$
$\widehat{\operatorname{FDP}}(t)_{\text{e-Ada-GAIF}}=\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}\frac{\mathbb{I}\{e_{j}\leq 1/\lambda\}}{1-\lambda}+\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}}{1+R({j-1})}\frac{\mathbb{I}\{e_{j}\leq 1/\lambda\}}{1-\lambda}.$
Then we require $\widehat{\operatorname{FDP}}_{\text{e-GAIF}}\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\text{e-Ada-GAIF}}\leq\alpha$ when constructing the testing levels. The corresponding testing levels are as follows:
$\alpha_{t}^{\text{e-GAIF}}=\omega_{t}\left(\alpha-\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}-\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}}{1+R({j-1})}\right)(R({t-1})+1),$ (21)
$\alpha_{t}^{\text{e-Ada-GAIF}}=\omega_{t}\left(\alpha(1-\lambda)-\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})\mathbb{I}\{e_{j}<1/\lambda\}}{1+R({j-1})}-\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}\mathbb{I}\{e_{j}<1/\lambda\}}{1+R({j-1})}\right)(R({t-1})+1),$ (22)
@@
The Generalized Alpha-Investing with Feedback procedure based on $e$-values is summarized in Algorithm 3. The $e$-GAIF and $e$-Ada-GAIF can control online FDR validly if the null $e$-values satisfy conditionally valid.
Algorithm 3 $e$-GAIF ($e$-GAIF and $e$-Ada-GAIF)
@@
3:   Update $\alpha_{t}=\alpha_{t}^{\text{e-GAIF}}$ in Eq.(21) (or $\alpha_{t}=\alpha_{t}^{\text{e-Ada-GAIF}}$ in Eq.(22))
@@
Theorem 6 (Online FDR control for e-GAIF and e-Ada-GAIF).
  • (a.)
@@
    where $\mathcal{F}_{t-1}=\sigma(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t-1}})$ is the sigma field generated from past rejections and feedback, then if the parameters $\{\alpha_{t}\}_{t\in\mathbb{N}}$ are selected such that $\widehat{\operatorname{FDP}}_{\operatorname{e-GAIF}}(t)\leq\alpha$, then we have